# R-Bot: An LLM-based Query Rewrite System

### v1

@@ 9. Conclusion
In this paper, we proposed an LLM-based query rewrite system. First, we prepared rewrite evidences from diverse sources, including rewrite rule specifications and rewrite Q&As. Next, we proposed a hybrid structure-semantics retrieval method to retrieve relevant rewrite evidences, based on which we generated rewrite recipes to instruct LLM for query rewrite. Then, we proposed a step-by-step LLM method, which iteratively utilized the retrieved Q&As and rewrite recipes to select and arrange rewrite rules with self-reflection. Experimental results demonstrated that R-Bot achieved remarkable improvements over existing query rewrite methods.

### v2

@@ 9. Conclusion
We proposed an LLM-based query rewrite system. First, we prepared rewrite evidences from diverse sources, including rewrite rule specifications and rewrite Q&As. Next, we proposed a hybrid structure-semantics retrieval method to retrieve relevant rewrite evidences, based on which we generated rewrite recipes to instruct LLM for query rewrite. Then, we proposed a step-by-step LLM method, which iteratively utilized the retrieved Q&As and rewrite recipes to select and arrange rewrite rules with self-reflection. Experimental results demonstrated that R-Bot achieved remarkable improvements over existing methods. Moreover, R-Bot deployed at Huawei and with real customers showed the effectiveness of R-Bot.
Acknowledgements.
This paper was supported by National Key R&D Program of China (2023YFB4503600), NSF of China (62525202, 62232009), Shenzhen Project (CJGJZD20230724093403007), Zhongguancun Lab, Huawei, and Beijing National Research Center for Information Science and Technology (BNRist). Guoliang Li is the corresponding author.